OpenStudy (anonymous):

tell weather the system has no solution, infintiely, many solution, or no solution 5x- 3y = 10 10x + 6y = 20

OpenStudy (tkhunny):

Multiply the first equation by 2. You should see something magic.

OpenStudy (anonymous):

i got 10x - 6y = 20 10x + 6y = 20

OpenStudy (anonymous):

If you add the equations (adding left to left and right to right), you will see the "y" term drop out. That means that you can solve for "x". After you get your "x" value, you can substitute that "x" value into either equation to get a "y" value. This is a system of 2 lines. If you get one solution, that means that they are not parallel and intersect at one point.

OpenStudy (anonymous):

u could show me please @tcarroll010

OpenStudy (anonymous):

I was looking for you to show me instead.

OpenStudy (anonymous):

so take the y out and add the x wish equal 20X and then add the two = set whish equal to 40 so i divide 40 out of 20 right @tcarroll010

OpenStudy (anonymous):

add left to left and right to right of the new set of equations.

OpenStudy (anonymous):

so i add 20 + 6+ 10 and so on

OpenStudy (anonymous):

by adding each side to each side, left to left and right to right, you just gather up all the terms on the left and simplify. You do the same for the right-hand side.

OpenStudy (anonymous):

i dont get u start it off for me

OpenStudy (anonymous):

start by identifying all the terms on the left sides in both equations, the NEW equations. Here's a big hint: there are 2 terms on the left-hand side of the first equation, and there are 2 terms on the left-hand side of the second equation. That's a total of 4 terms. You put all 4 terms on the same left-hand side in a new combined equation. You do the same for the right-hand side. Then you simplify. That's the biggest hint I can give you without just giving away an answer.

OpenStudy (anonymous):

i get 20x + 0y = 40

OpenStudy (anonymous):

And 0y is 0. So, you solved for x by dividing each side by 20.

OpenStudy (anonymous):

i get o for the righ side and i get 10 for the other side

OpenStudy (anonymous):

You are working with: 20x = 40 You divide both sides by 20 to solve for "x".

OpenStudy (anonymous):

x equal 2 so it is a one solution

OpenStudy (anonymous):

Yes! Good job!

OpenStudy (anonymous):

thank u very much @tcarroll010